# Supernovae in 2023 (review): breakthroughs by late observations

### v3

@@ 3.1 Neutrino-driven versus jet-driven explosions
Some CCSNe require a central engine to operate after the explosion itself (e.g., Matsumoto, Metzger, & Goldberg 2024). Many studies attribute the extra energy to a rapidly rotating magnetized NS, i.e., a magnetar (e.g., Acharyya et al. 2023; Bai et al. 2023; Chen et al. 2023c; Gkini et al. 2023; Hu et al. 2023b; Omand & Sarin 2023; Prasanna et al. 2023; Tinyanont et al. 2023; Wang et al. 2023). As I pointed out in the past, the formation process of an energetic magnetar is accompanied by the launching of even more energetic jets, during and/or after the explosion. The majority of magnetar studies continue to ignore the role of jets. The problems of fitting the lightcurves of superluminous CCSNe with magnetars continue with 2023 papers. In Soker (2022) I analyzed the paper by Chen et al. (2023d) who fitted the lightcurves of 70 hydrogen-poor superluminous CCSNe with magnetars, about a third of these have an initial spin period of $P_{\rm mag}\leq 0.002{~{}\rm s}$. However, such rapidly rotating magnetars are likely to be spun up during the last phase of their formation by an accretion disk/belt that launches jets. I argued there that even the slower magnetars in their fitting most likely launched jets at the explosion. In seven out of the 14 CCSNe they fit with ejecta-CSM interaction, the explosion energy is much larger than what the delayed neutrino explosion mechanism can supply, implying a jet-driven explosion. The claim that supernova studies that include energetic magnetars cannot ignore jets continues to hold in 2023. The same arguments hold for the magnetars with $P_{\rm mag}\leq 0.002{~{}\rm s}$ that West et al. (2023) and Poidevin et al. (2023) find to fit the hydrogen-poor superluminous SN 2020qlb and SN 2021fpl, respectively, for the number of short-period magnetars that Dong et al. (2023c) argue to power bumpy lightcurves, for those magnetars with $P_{\rm mag}\leq 0.002{~{}\rm s}$ that Gomez et al. (2023) and Hinkle, Shappee, & Tucker (2023) fit to hydrogen-poor superluminous CCSNe, and to the magnetar that Fiore et al. (2023) fit to SN 2019neq.
@@
The delayed neutrino explosion mechanism cannot account for energetic CCSNe. Hiramatsu et al. (2023a) find the explosion energy of the type IIn SN 2021qqp to be $E_{\rm exp}\simeq(3-10)\times 10^{51}{~{}\rm erg}$, and Smith et al. (2023) estimate $E_{\rm exp}\simeq(5-10)\times 10^{51}{~{}\rm erg}$ for type IIn SN 2015da. The delayed neutrino explosion mechanism cannot account for such explosion energies, while the JJEM can naturally explain such explosion energies (for a discussion from 2023 see, e.g., Shishkin & Soker 2023). Salmaso et al. (2023) deduced that the explosion energy of SN 2020faa is $\simeq 4\times 10^{51}{~{}\rm erg}$, and discussed the possibility that this peculiar SN II was powered mainly by jets. Likewise, Nagao et al. (2023a) find the explosion energy of the type Icn SN 2021ckj to be $\simeq 4\times 10^{51}{~{}\rm erg}$, an energy the neutrino delayed mechanism cannot account for. Pumo et al. (2023) specifically argue that some of their modeled CCSNe that have high explosion energies challenge the neutrino-driven explosion mechanism. Schweyer et al. (2023) estimate the kinetic energy of SN 2019odp, a type Ib CCSN, to be $E_{\rm exp}\simeq 6\times 10^{51}{~{}\rm erg}$ and note that it is above what the delayed neutrino explosion mechanism can supply. Karamehmetoglu et al. (2023) model several SNe Ibc and find their kinetic energy to be $>3\times 10^{51}{~{}\rm erg}$, more than what the delayed neutrino explosion mechanism can account for. Again, the JJEM easily accounts for such explosion energies. Ergon et al. (2023) discuss the possibility that jets can improve their fitting of the lightcurve of the type IIb SN 2020acat.
@@
Another unsolved problem of the delayed neutrino explosion mechanism is that it seems that it cannot produce the observed mass of ${}^{56}$i in stripped-envelope CCSNe (Sawada & Suwa, 2023). Imasheva, Janka, & Weiss (2023), on the other hand, argue that with proper treatment of the collapsing core, they can produce more ${}^{56}$i, at least for the model appropriate for SN 1987A.
@@
Studies in 2023 continue the exploration of gravitational waves from CCSNe and their detectability (e.g., Bruel et al. 2023; Bugli et al. 2023; Casallas Lagos et al. 2023; Dálya et al. 2023; Heston et al. 2023; Hsieh et al. 2023; Iess et al. 2023; Jakobus et al. 2023; Mezzacappa et al. 2023; Mitra et al. 2023; Pajkos et al. 2023; Pastor-Marcos et al. 2023; Shibagaki et al. 2023; Szczepańczyk et al. 2023; Vartanyan et al. 2023; Villegas et al. 2023; Wolfe et al. 2023; Yuan et al. 2023; Nunes et al. 2024; Zha, Eggenberger Andersen, & O’Connor 2024a), including from two cocoons (lobes) that are inflated by fixed-axis energetic jets related to CCSNe (Gottlieb et al., 2023) and by several jet-inflated pairs of lobes (cocoons) in the JJEM (Soker, 2023c). The detection of neutrinos from CCSNe continued to be a topic of interest in 2023 (e.g., Abbasi et al. 2023; Akaho, Nagakura, & Foglizzo 2023; Bendahman et al. 2023; Drago et al. 2023; Fiorillo et al. 2023; Lin et al. 2023b; Mukhopadhyay & Sen 2023; Pagliaroli & Ternes 2024; Prasanna, Thompson, & Hirata 2023; Vartanyan & Burrows 2023; Ashida 2024; Powell et al. 2023; Saez et al. 2024). The detection of gravitational waves or neutrinos is not at a stage that allows the support of one explosion mechanism or the other (for the study of CCSN diagnostics, see, e.g., Fryer et al. 2023). Therefore, I will not touch on these topics. I will also not touch the subject of axion instability supernova (e.g., Mori et al. 2023).
@@
The presence of CSM around CCSN progenitors and the interaction of the CCSN ejecta with this CSM were major subjects in the 2023 supernova literature (e.g., Akashi, Michaelis, & Soker 2023; Andrews et al. 2023; Ben-Ami et al. 2023; Bietenholz et al. 2023; Bilinski et al. 2023; Bruch et al. 2023; Chugai & Utrobin 2023; Das et al. 2023; Davis et al. 2023; Dessart & Jacobson-Galán 2023; Dickinson et al. 2023; Ercolino et al. 2023; Hiramatsu et al. 2023a; Ishii et al. 2023; Khatami & Kasen 2023; Koo et al. 2023a; Lin et al. 2023; Maeda et al. 2023a, c; Matsuoka & Sawada 2023; Mauerhan et al. 2023; Moran et al. 2023; Moriya 2023; Pearson et al. 2023a; Pellegrino et al. 2023; Pessi et al. 2023a; Petruk et al. 2023; Pursiainen et al. 2023a; Salmaso et al. 2023; Sarmah et al. 2023; Shrestha et al. 2023; Takei et al. 2023; Tsuna, Murase, & Moriya 2023; Tsuna & Takei 2023; Brennan et al. 2024a; Ferrari et al. 2024; Jacobson-Galán et al. 2024; Murase 2024; Meza Retamal et al. 2024; Murai et al. 2024; Sfaradi et al. 2024), including the role of binary interaction in expelling pre-explosion CSM (e.g., Zenati et al. 2024). [^1] Soker (2013) suggested (based on earlier binary models of similar types of transients) that a sudden pre-CCSN swelling may trigger asymmetric mass loss in binary systems. Five months later Smith & Arnett (2014) repeated this suggestion. Later papers (e.g., Smith 2017; Smith et al. 2023; Bilinski et al. 2023; Mauerhan et al. 2023; Smith et al. 2023) wrongly and consistently have been attributing this suggestion to Smith & Arnett (2014). The ejecta-CSM interaction has become a major topic in relation to the type II SN 2023ixf (e.g., Bersten et al. 2023; Chandra et al. 2023; Guetta et al. 2023; Hiramatsu et al. 2023b; Hosseinzadeh et al. 2023a; Koenig 2023; Liu et al. 2023a; Martinez et al. 2023; Niu et al. 2023; Pledger & Shara 2023; Qin et al. 2023; Sarmah 2023; Sgro et al. 2023; Teja et al. 2023; Van Dyk et al. 2023b; Vasylyev et al. 2023b; Xiang et al. 2023a; Yamanaka, Fujii, & Nagayama 2023; Zimmerman et al. 2023).
@@
Studies found that the ejecta of SN 2023ixf interacted with a CSM that was extended up to a radius of $R_{\rm CSM}\simeq 20-50{~{}\rm AU}$ around SN 2023ixf (e.g., Berger et al. 2023; Bostroem et al. 2023b; Grefenstette et al. 2023; Jacobson-Galán et al. 2023; Kilpatrick et al. 2023a; Teja et al. 2023; Smith et al. 2023; Zhang et al. 2023a; Li et al. 2023). In principle, a close CSM might result from an enhanced mass loss rate shortly before the explosion, most likely accompanied by an outburst (eruption; e.g., Tsuna, Takei, & Shigeyama 2023; Tsuna et al. 2024) that might be triggered by activity in the core (e.g. Cohen & Soker 2024). However, several studies find no evidence for pre-explosion outbursts in the years before the explosion of SN 2023ixf (e.g., Dong et al. 2023a; Flinner et al. 2023; Neustadt, Kochanek, & Smith 2023; Jencson et al. 2023; Soraisam et al. 2023; Ransome et al. 2024). This challenges any model where the massive CSM is attributed to impulsive mass ejection years before the explosion of SN 2023ixf. The progenitor of SN 2023ixf did have large amplitude pulsations (e.g., Kilpatrick et al. 2023a; Soraisam et al. 2023). These pre-explosion properties led to the suggestion that instead of a pre-explosion impulsive ejection of a massive CSM, the progenitor of SN 2023ixf maintained an extended zone above its photosphere of rising and falling dense clumps, i.e., the effervescent zone model (Soker, 2023a).
@@
Obergaulinger & Reichert (2023) study nucleosynthesis in fixed-axis jets that power some energetic CCSNe, and find that they have enhanced synthesis of iron-group elements. Reichert et al. (2024) find that when magnetic fields are included in the fixed-axis simulations, nucleosynthesis can yield up to the third peak of the r-process (but see Zha, Müller, & Powell 2024b). Leung, Nomoto, & Suzuki (2023) and Leung & Nomoto (2024) study nucleosynthesis in CCSNe driven by fixed-axis jets, including collapsar where the remnant is a black hole. Wanajo (2023) studies nucleosynthesis in neutrino-heated ejecta and neutrino-driven winds (see also Prasanna, Coleman, & Thompson (2024)), and Friedland, Mukhopadhyay, & Patwardhan (2024) and Psaltis et al. (2023) study the $\nu p$-process in neutrino-driven outflows in CCSNe. These papers do not refer to jets. In general, r-process nucleosynthesis in CCSNe was not a major topic in 2023. Anand et al. (2023), for example, study the lightcurves of broad-line stripped-envelope CCSNe and find no evidence for r-process nucleosynthesis. Patel et al. (2024) study the effect of r-process nucleosynthesis on the lightcurve of hydrogen-rich CCSNe. Nucleosynthesis can be used to study CCSNe. Hoppe et al. (2023) and Liu et al. (2024) present interesting studies of using solar-system isotopes ratio to constrain the stellar properties at the explosion.
@@
Several papers in 2023 (e.g., Reichert et al. 2023; Volpato et al. 2023) study the rare group of ‘magnetorotational supernovae’, which are driven by fixed-axis jets due to rapid pre-collapse core rotation. While in the delayed neutrino explosion mechanism these rare CCSNe form a separate group (e.g., Fujibayashi et al. 2023a; Maeda, Suzuki, & Izzo 2023), in the JJEM they are just the rapidly rotating tail of CCSNe with slow pre-collapse core rotation (e.g., Soker 2023b). In the fixed-axis case the rapid pre-collapse rotation of the core forces the accretion disk around the newly born NS (and in some cases later around the black hole) to maintain a constant axis, and so are the jets that the accretion disk launches. This leads to a bipolar explosion (e.g., Fang et al. 2023). Vasylyev et al. (2024) measure high continuum linear polarization in the type IIP supernova 2021yja, and discuss the possibilities that it results from CSM interaction or fixed-axis jet explosion. Polarization measurements can teach a lot about the explosion mechanism (e.g., Maund 2023). Nagao et al. (2023b) conclude their spectropolarimetry study of SNe II by stating that “the emergence of a global aspherical structure, e.g., a jet-like structure, might be the key ingredient in the explosion mechanism to produce an energetic SN”. This is compatible with the expectation of the JJEM.
However, such a rapid rotation is rare. On the other extreme, the core rotation is zero, and the angular momentum of the intermittent disks that launch the jets is stochastic such that the jets are fully jittering. The stochastic angular momentum direction results from seed angular momentum perturbations, by the pre-collapse core convection, that are amplified by instabilities above the newly born NS. In the JJEM there is a continuum from CCSNe driven by fully jittering jets, which result from progenitors with very slowly rotating pre-collapse cores, to CCSNe driven by fixed-axis jets that result from rapidly rotating pre-collapse cores. In the frame of the JJEM, this behavior can explain the mass gap between NSs and black holes (Soker, 2023b).

### v4

@@ 3.1 Neutrino-driven versus jet-driven explosions
Some CCSNe require a central engine to operate after the explosion itself (e.g., Matsumoto, Metzger, & Goldberg 2024). Many studies attribute the extra energy to a rapidly rotating magnetized NS, i.e., a magnetar (e.g., Acharyya et al. 2023; Bai et al. 2023; Chen et al. 2023c; Gkini et al. 2023; Hu et al. 2023b; Omand & Sarin 2024; Prasanna et al. 2023; Tinyanont et al. 2023; Wang et al. 2023c). As I pointed out in the past, the formation process of an energetic magnetar is accompanied by the launching of even more energetic jets during and/or after the explosion. The majority of magnetar studies continue to ignore the role of jets. The problems of fitting the lightcurves of superluminous CCSNe with magnetars continue with 2023 papers. In Soker (2022c), I analyzed the paper by Chen et al. (2023d), who fitted the lightcurves of 70 hydrogen-poor superluminous CCSNe with magnetars; about a third of these have an initial spin period of $P_{\rm mag}\leq 0.002{~{}\rm s}$. However, such rapidly rotating magnetars will likely be spun up during the last phase of their formation by an accretion disk/belt that launches jets. I argued that even the slower magnetars in their fitting most likely launched jets at the explosion. In seven out of the 14 CCSNe they fit with ejecta-CSM interaction, the explosion energy is much larger than what the delayed neutrino explosion mechanism can supply, implying a jet-driven explosion. The claim that supernova studies that include energetic magnetars cannot ignore jets continues to hold in 2023. The same arguments hold for the magnetars with $P_{\rm mag}\leq 0.002{~{}\rm s}$ that West et al. (2023) and Poidevin et al. (2023) find to fit the hydrogen-poor superluminous SN 2020qlb and SN 2021fpl, respectively, for the number of short-period magnetars that Dong et al. (2023c) argue to power bumpy lightcurves, for those magnetars with $P_{\rm mag}\leq 0.002{~{}\rm s}$ that Gomez et al. (2023) and Hinkle, Shappee, & Tucker (2023) fit to hydrogen-poor superluminous CCSNe, and to the magnetar that Fiore et al. (2024) fit to SN 2019neq.
@@
The delayed neutrino explosion mechanism cannot account for energetic CCSNe. Hiramatsu et al. (2023a) find the explosion energy of the type IIn SN 2021qqp to be $E_{\rm exp}\simeq(3-10)\times 10^{51}{~{}\rm erg}$, and Smith et al. (2023) estimate $E_{\rm exp}\simeq(5-10)\times 10^{51}{~{}\rm erg}$ for type IIn SN 2015da. The delayed neutrino explosion mechanism cannot account for such explosion energies, while the JJEM can naturally explain such explosion energies (for a discussion from 2023, see, e.g., Shishkin & Soker 2023). Salmaso et al. (2023) deduced that the explosion energy of SN 2020faa is $\simeq 4\times 10^{51}{~{}\rm erg}$, and discussed the possibility that this peculiar SN II was powered mainly by jets. Likewise, Nagao et al. (2023) find the explosion energy of the type Icn SN 2021ckj to be $\simeq 4\times 10^{51}{~{}\rm erg}$, an energy the neutrino delayed mechanism cannot account for. Pumo et al. (2023) specifically argue that some of their modeled CCSNe with high explosion energies challenge the neutrino-driven explosion mechanism. Schweyer et al. (2023) estimate the kinetic energy of SN 2019odp, a type Ib CCSN, to be $E_{\rm exp}\simeq 6\times 10^{51}{~{}\rm erg}$ and note that it is above what the delayed neutrino explosion mechanism can supply. Karamehmetoglu et al. (2023) model several SNe Ibc and find their kinetic energy to be $>3\times 10^{51}{~{}\rm erg}$, more than what the delayed neutrino explosion mechanism can account for. Again, the JJEM easily accounts for such explosion energies. Ergon et al. (2023) discuss the possibility that jets can improve their fitting of the lightcurve of the type IIb SN 2020acat.
@@
Another unsolved problem of the delayed neutrino explosion mechanism is that it seems it cannot produce the observed mass of 56Ni in stripped-envelope CCSNe (Sawada & Suwa, 2023). Imasheva, Janka, & Weiss (2023), on the other hand, argue that with proper treatment of the collapsing core, they can produce more 56Ni, at least for the model appropriate for SN 1987A.
@@
Studies in 2023 continue the exploration of gravitational waves from CCSNe and their detectability (e.g., Bruel et al. 2023; Bugli et al. 2023; Casallas Lagos et al. 2023; Dálya et al. 2023; Heston et al. 2023; Hsieh et al. 2023; Iess et al. 2023; Jakobus et al. 2023; Mezzacappa et al. 2023; Mitra et al. 2023; Pajkos et al. 2023; Pastor-Marcos et al. 2023; Shibagaki et al. 2023; Szczepańczyk et al. 2023; Vartanyan et al. 2023; Villegas et al. 2023; Wolfe et al. 2023; Yuan et al. 2023; Nunes et al. 2024; Zha, Eggenberger Andersen, & O’Connor 2024a), including from two cocoons (lobes) that are inflated by fixed-axis energetic jets related to CCSNe (Gottlieb et al., 2023) and by several jet-inflated pairs of lobes (cocoons) in the JJEM (Soker, 2023c). The detection of neutrinos from CCSNe continued to be a topic of interest in 2023 (e.g., Abbasi et al. 2023; Akaho, Nagakura, & Foglizzo 2024; Bendahman et al. 2024; Drago et al. 2023; Fiorillo et al. 2023; Lin et al. 2023b; Mukhopadhyay & Sen 2024; Pagliaroli & Ternes 2024; Prasanna, Thompson, & Hirata 2024; Vartanyan & Burrows 2023; Ashida 2024; Powell et al. 2024; Saez et al. 2024). The detection of gravitational waves or neutrinos is not at a stage that allows the support of one explosion mechanism or the other (for the study of CCSN diagnostics, see, e.g., Fryer et al. 2023). Therefore, I will not touch on these topics. I will also not touch the subject of axion instability supernova (e.g., Mori et al. 2023).
@@
The presence of CSM around CCSN progenitors and the interaction of the CCSN ejecta with this CSM were major subjects in the 2023 supernova literature (e.g., Akashi, Michaelis, & Soker 2023; Andrews et al. 2023; Ben-Ami et al. 2023; Bietenholz et al. 2023; Bilinski et al. 2024; Bruch et al. 2023; Chugai & Utrobin 2023; Davis et al. 2023; Dessart & Jacobson-Galán 2023; Dickinson et al. 2023; Ercolino et al. 2023; Hiramatsu et al. 2023a; Ishii et al. 2024; Khatami & Kasen 2023; Koo et al. 2023a; Lin et al. 2023; Maeda et al. 2023a, c; Matsuoka & Sawada 2024; Mauerhan et al. 2024; Moran et al. 2023; Moriya 2023; Pearson et al. 2023; Pellegrino et al. 2023; Pessi et al. 2023a; Petruk et al. 2023; Pursiainen et al. 2023a; Salmaso et al. 2023; Sarmah et al. 2023; Shrestha et al. 2024; Takei et al. 2024; Tsuna, Murase, & Moriya 2023; Tsuna & Takei 2023; Brennan et al. 2024a; Das et al. 2024b; Ferrari et al. 2024; Jacobson-Galán et al. 2024; Murase 2024; Meza Retamal et al. 2024; Murai et al. 2024; Sfaradi et al. 2024), including the role of binary interaction in expelling pre-explosion CSM (e.g., Zenati et al. 2024). [^1] Soker (2013) suggested (based on earlier binary models of similar types of transients) that a sudden pre-CCSN swelling may trigger asymmetric mass loss in binary systems. Five months later, Smith & Arnett (2014) repeated this suggestion. Later papers (e.g., Smith 2017; Smith et al. 2023; Bilinski et al. 2024; Mauerhan et al. 2024; Smith et al. 2023) wrongly and consistently have been attributing this suggestion to Smith & Arnett (2014). The ejecta-CSM interaction has become a major topic in relation to the type II SN 2023ixf (e.g., Bersten et al. 2024; Chandra et al. 2024; Guetta et al. 2023; Hiramatsu et al. 2023b; Hosseinzadeh et al. 2023a; Koenig 2023; Liu et al. 2023a; Martinez et al. 2024; Niu et al. 2023; Pledger & Shara 2023; Qin et al. 2023; Sarmah 2023; Sgro et al. 2023; Teja et al. 2023; Van Dyk et al. 2023b; Vasylyev et al. 2023b; Xiang et al. 2024; Yamanaka, Fujii, & Nagayama 2023; Zimmerman et al. 2023).
@@
Studies found that the ejecta of SN 2023ixf interacted with a CSM that was extended up to a radius of $R_{\rm CSM}\simeq 20-50{~{}\rm AU}$ around SN 2023ixf (e.g., Berger et al. 2023; Bostroem et al. 2023b; Grefenstette et al. 2023; Jacobson-Galán et al. 2023; Kilpatrick et al. 2023a; Teja et al. 2023; Smith et al. 2023; Zhang et al. 2023a; Li et al. 2023). In principle, a close CSM might result from an enhanced mass loss rate shortly before the explosion, most likely accompanied by an outburst (eruption; e.g., Tsuna, Takei, & Shigeyama 2023; Tsuna et al. 2024) that might be triggered by activity in the core (e.g. Cohen & Soker 2024). However, several studies find no evidence for pre-explosion outbursts in the years before the explosion of SN 2023ixf (e.g., Dong et al. 2023a; Flinner et al. 2023; Neustadt, Kochanek, & Smith 2024; Jencson et al. 2023; Soraisam et al. 2023; Ransome et al. 2024). This challenges any model where the massive CSM is attributed to impulsive mass ejection years before the explosion of SN 2023ixf. The progenitor of SN 2023ixf did have large amplitude pulsations (e.g., Kilpatrick et al. 2023a; Soraisam et al. 2023). These pre-explosion properties led to the suggestion that instead of a pre-explosion impulsive ejection of a massive CSM, the progenitor of SN 2023ixf maintained an extended zone above its photosphere of rising and falling dense clumps, i.e., the effervescent zone model (Soker, 2023a).
@@
Obergaulinger & Reichert (2023) study nucleosynthesis in fixed-axis jets that power some energetic CCSNe, and find that they have enhanced synthesis of iron-group elements. Reichert et al. (2024) find that when magnetic fields are included in the fixed-axis simulations, nucleosynthesis can yield up to the third peak of the r-process (but see Zha, Müller, & Powell 2024b). Leung, Nomoto, & Suzuki (2023) and Leung & Nomoto (2024) study nucleosynthesis in CCSNe driven by fixed-axis jets, including collapsar where the remnant is a black hole. Wanajo (2023) studies nucleosynthesis in neutrino-heated ejecta and neutrino-driven winds (see also Prasanna, Coleman, & Thompson 2024), and Friedland, Mukhopadhyay, & Patwardhan (2024) and Psaltis et al. (2023) study the $\nu p$-process in neutrino-driven outflows in CCSNe. These papers do not refer to jets. In general, r-process nucleosynthesis in CCSNe was not a major topic in 2023. Anand et al. (2024), for example, study the lightcurves of broad-line stripped-envelope CCSNe and find no evidence for r-process nucleosynthesis. Patel et al. (2024) study the effect of r-process nucleosynthesis on the lightcurve of hydrogen-rich CCSNe. Nucleosynthesis can be used to study CCSNe. Hoppe et al. (2023) and Liu et al. (2024) present interesting studies of using solar-system isotopes ratio to constrain the stellar properties at the explosion.
@@
Several papers in 2023 (e.g., Reichert et al. 2023; Volpato et al. 2023) study the rare group of ‘magnetorotational supernovae’, which are driven by fixed-axis jets due to rapid pre-collapse core rotation. While in the delayed neutrino explosion mechanism, these rare CCSNe form a separate group (e.g., Fujibayashi et al. 2023; Maeda, Suzuki, & Izzo 2023), in the JJEM, they are just the rapidly rotating tail of CCSNe with slow pre-collapse core rotation (e.g., Soker 2023b). In the fixed-axis case, the rapid pre-collapse rotation of the core forces the accretion disk around the newly born NS (and in some cases later around the black hole) to maintain a constant axis, and so are the jets that the accretion disk launches. This leads to a bipolar explosion (e.g., Fang et al. 2024). Vasylyev et al. (2024) measure high continuum linear polarization in the type IIP supernova 2021yja and discuss the possibilities that it results from CSM interaction or fixed-axis jet explosion. Polarization measurements can teach much about the explosion mechanism (e.g., Maund 2024). Nagao et al. (2024) conclude their spectropolarimetry study of SNe II by stating that “the emergence of a global aspherical structure, e.g., a jet-like structure, might be the key ingredient in the explosion mechanism to produce an energetic SN”. This is compatible with the expectation of the JJEM.
However, such a rapid rotation is rare. On the other extreme, the core rotation is zero, and the angular momentum of the intermittent disks that launch the jets is stochastic, such that the jets are fully jittering. The stochastic angular momentum direction results from seed angular momentum perturbations by the pre-collapse core convection that are amplified by instabilities above the newly born NS. In the JJEM, there is a continuum from CCSNe driven by fully jittering jets, which result from progenitors with very slowly rotating pre-collapse cores to CCSNe driven by fixed-axis jets that result from rapidly rotating pre-collapse cores. In the frame of the JJEM, this behavior can explain the mass gap between NSs and black holes (Soker, 2023b).